# Flipped rotating axion non-minimally coupled to gravity: Baryogenesis and Dark Matter

### v1

@@ VII Fragmentation of axion
Our setup can lead to the enhancement of axion fluctuations, which can lead to a significant loss of kinetic energy of the rotating axion, eventually stopping the motion of the axion. This can happen through parametric resonance of certain values of wave numbers. This phenomenon known as axion fragmentation was first studied in details in Ref. Fonseca et al. (2020) and further explored in Refs. Eröncel et al. (2022); Eröncel and Servant (2023). In Ref. Eröncel et al. (2022), it was shown that the axion can be fragmented even before getting trapped, if the mass is much larger than the Hubble rate during trapping, i.e., $m_{\phi}/H_{\rm ufr}\gg\mathcal{O}(1)$. Basically, this means that when $H$ reaches $m_{\phi}$, the axion still has larger kinetic energy than the barrier, thus delaying the conventional oscillation time. Recall that in our case the trapping potential given by Eq. (3) is different from the initial potential (Eq. (7)) over which the axion starts to roll, and we have $m_{\phi}\simeq H_{\rm ufr}$ during trapping, at least for the case when $H_{\rm fr}>H_{\rm ufr}$ (Eq. (34)). However, during the initial rotation in the kination domination, parametric resonance may be important, as it can stop the rotation of our axion. We investigate this danger below.
@@
where we used Eq. (19) for the initial velocity $\dot{\theta}_{m}$ and 12(${}^{2}\simeq\,$5.3. Note that violating the above condition doesn’t necessarily implies that the axion would be fragmented. To see if the modes would grow, one has to integrate the growth factor within the exponent over the time that the modes stay inside the instability band Fonseca et al. (2020); Eröncel et al. (2022), which can be very small. In this work, we don’t derive a full analytical expression for this growth but rather solve Eq. (52) numerically along with the background homogeneous equation [^9] The equation of motion for $\theta$ is given by $\ddot{\theta}+3H\dot{\theta}+\left[\frac{k^{2}}{a^{2}}+V^{\prime}(\phi)\right]\theta=0$, where $V^{\prime}(\phi)=\frac{3\xi m_{P}^{2}H^{2}}{f^{2}}\sin(\frac{\phi}{f})$. for $\theta$, to see if our condition in Eq. (56) holds [^10] We consider adiabatic initial conditions for Eq. (52), sourced by the curvature perturbations Mukhanov (2005); Eröncel et al. (2022); Eröncel and Servant (2023)..
@@
Therefore, the lower part of the range in Eq. (9) is safe against fragmentation of the rotating axion condensate, for which we can have successfull baryogenesis as well, as discussed in Section III. In fact, in the next section we argue that it may be preferable for $\xi$ to be not much larger than $(f/m_{P})^{2}$.
We leave a detailed study of axion fragmentation including improved analytical results and numerical solutions including backreaction in such a scenario for future studies.
@@ VIII The Kibble issue
Therefore, we expect the axion to roll only in one direction (e.g. clockwise) in the observable Universe, which evades the concerns of the Kibble problem. The price to pay, if this axion is also the dark matter, is that we have to make sure that excessive isocurvature perturbations are avoided. [^12] This is quite feasible for a rotating axion, e.g. see Ref. Chung and Tadepalli (2024).
Another important point is that $\xi$ has to gradually increase during kination if the axion is to roll at all, because the latter requires $||V_{\rm eff}^{\prime\prime}||>H^{2}$. For this to happen $\xi$ needs to grow by a couple of orders of magnitude, if we are close to the upper bound in Eq. (58). One possibility is to consider a running $\xi(\sigma)$ as Bezrukov and Shaposhnikov (2014); Hamada et al. (2015); Ezquiaga et al. (2018); Drees and Xu (2021); Cheong et al. (2021); Ghoshal et al. (2024)
@@
i.e. near the edge of both ranges in Eqs. (9) and (58). It is such a value that was assumed in the examples at the end of Sec. IV, while, as shown in Sec. VII, it is also safe from fragmentation of the axion condensate.
Combining Eqs. (24) and (60), it is straightforward to find
@@
where we used $Y_{B}\sim 10^{-10}$. Thus, if for example $T_{\rm reh}\sim 10^{10}\,$eV we find $T_{\rm B-L}\sim 10^{9}\,$eV. Considering the BBN bound $T_{\rm reh}\gtrsim 10^{7}$ GeV, we get $T_{\rm B-L}\gtrsim 3\times 10^{7}$ GeV. Thus, for $z_{\rm fo}$ around $\mathcal{O}(10)$, as discussed in Section VI, this gives a lower bound on the RHN mass as $M_{N}\gtrsim 3\times 10^{8}$ GeV.
@@ IX Conclusion
In this work, we investigated an attractive way to generate rotation for an axion-like particle, if it is non-minimally coupled to gravity, respecting the shift-symmetry. Such a rotation is generated when the effective potential of the axion flips at the end of inflation, providing the axion enough kinetic energy to surpass the potential barrier. This becomes possible if inflation is followed by a kination-dominated epoch ($w=1$), when the axion potential barrier redshifts in the same way as its kinetic energy, i.e. as $a^{-6}$. We find that for the axion to roll and remain sub-dominant, the value of $\xi$ is constrained to be in the range given by Eq. (9). We utilize such an axion rotation to generate the baryon asymmetry through spontaneous baryogenesis. The final baryon asymmetry, given by Eq. (24), is found to depend on the initial kinetic energy of the axion determined crucially by the non-minimal coupling $\xi$ and the decay constant $f$, and its subsequent redshift which depends on the reheating temperature $T_{\rm reh}$. Additionally, the asymmetry is determined by the decoupling temperature of the baryon or lepton number violating interaction $T_{\rm B-L}$, which depends on the details of the particle physics setup. The axion continues to rotate even after reheating until the kinetic energy becomes comparable to the height of the bare mass potential (Eq. (3)), determined by $M$, after which the coherent oscillation of the axion can behave as dark matter, relating the axion decay constant to its mass through Eq. (33).
The presence of the kination-dominated epoch leads to a substantial enhancement of the stochastic gravitational wave background from inflation. The blue-tilt of such a GW spectrum is determined by the reheating temperature. In order not to overproduce GW during BBN, the reheating temperature is constrained to be greater than $\mathcal{O}(10^{7})\,$GeV (Eq. (47)), which substantially constrains the viable parameter space.
@@
We show a concrete example of Majoron in the well-motivated Type I seesaw model, where our framework can be utilized for the cogenesis of baryon asymmetry and dark matter, with the viable parameter space shown in Fig. 4. It is important to note that in our mechanism of flipped axion rotation, although the same axion field is responsible for cogenesis, the potential (Eq. (7)) playing role in baryogenesis during rotation differs from that (Eq. (3)) generating the dark matter abundance during oscillation. Thus, the axion mass $m_{\phi}$ is unconstrained from the observed baryon asymmetry (cf. Eq. (24)), which opens up a larger parameter space unlike in other axion rotation scenarios. In our example, the Majoron mass is generally predicted to take any value in the range below sub-eVs. On the other hand, the decoupling temperature $T_{\rm B-L}$ is tightly constrained by bounds from GW overproduction and issues related to fragmentation and Kibble problem discussed earlier. As a result, in our example, RHN mass which determines $T_{\rm B-L}$ is constrained to be larger than around $10^{8}$ GeV (see discussion below Eq. (61)). While considering $m_{\nu}\sim 0.05$ eV and the Yukawa couplings to obey the perturbativity bound, $M_{N}$ and $T_{\rm max}$ can be upto the GUT-scale.
Our scenario is very predictive and favors specific values of the non-minimal coupling $\xi$, which for instance, considering GUT-scale decay constant $f\sim 10^{-2}\leavevmode\nobreak\ m_{P}$ is around $10^{-4}$. These predictions can be indirectly tested by GW experiments like BBO, UDECIGO and resonant cavities, which can be sensitive to the blue tilt (see Fig. 3). Morever, the predicted values of $\Delta N_{\rm eff}$ coming from the GW background can be within the sensitivities of several experiments such as CMB-S4, CMB-HD, EUCLID (cf. Table 1, Figs. 3, 4). Our idea might also have interesting phenomenology for other particle physics model including the QCD axion, which we leave for future exploration.
@@ Acknowledgements
CC is supported by National Natural Science Foundation of China (No. 12433002) and Start-up Funds for Doctoral Talents of Jiangsu University of Science and Technology. SJD is supported by IBS under the project code IBS-R018-D1. KD is supported (in part) by the STFC consolidated Grant: ST/X000621/1. Authors thank Pawel Kozow for discussion. CC thanks the supports from The Asia Pacific Center for Theoretical Physics, The Center for Theoretical Physics of the Universe at Institute for Basic Science during his visits.

### v2

@@ VII Axion Fragmentation
Our setup can lead to the enhancement of axion fluctuations, which can lead to a significant loss of kinetic energy of the rotating axion, eventually stopping the variation of the axion. This can happen through parametric resonance of certain values of wave numbers. This phenomenon known as axion fragmentation was first studied in detail in Ref. Fonseca et al. (2020) and further explored in Refs. Eröncel et al. (2022); Eröncel and Servant (2023). In Ref. Eröncel et al. (2022), it was shown that the axion can be fragmented even before getting trapped, if the mass is much larger than the Hubble rate during trapping, i.e., $m_{\phi}/H_{\rm ufr}\gg\mathcal{O}(1)$. Basically, this means that when $H$ reaches $m_{\phi}$, the axion still has larger kinetic energy than the barrier, thus delaying the conventional oscillation time. Recall that in our case the trapping potential given by Eq. (3) is different from the initial potential (Eq. (7)) over which the axion starts to roll, and we have $m_{\phi}\simeq H_{\rm ufr}$ during trapping, at least for the case when $H_{\rm fr}>H_{\rm ufr}$ (subject to Eq. (36)). However, during the initial rotation in the kination domination, parametric resonance may be important, as it can stop the rotation of our axion. We investigate this danger below.
@@
[…]
@@
An upper bound on $\xi$ close to our numerical results can be derived as follows Eröncel et al. (2022). During the amplification, the velocity and the scale factor can be written as
$\dot{\theta}(t)=\dot{\theta}(t_{k})+\ddot{\theta}(t_{k})(t-t_{k})\,\,$,\,\,
$a(t)=a(t_{k})[1+H(t_{k})(t-t_{k})]\,$, (61)
where the subscript ‘$k$’ indicates the time when the mode is at the center of the instability band given by Eq. (58). Here, we consider $\ddot{\theta}$ and $H$ to be constant during the amplification. Using the above in Eq. (57), the total amplification time, i.e. the time spent inside the instability band, is found to be
$t_{\rm amp}=2|\Delta t|\simeq\frac{2\times 3\xi m_{P}^{2}H^{2}}{f^{2}\dot{\theta}(t_{k})|H(t_{k})\dot{\theta}(t_{k})+\ddot{\theta}(t_{k})|}\,.$ (62)
where $\Delta t(-\Delta t)$ corresponds to the time taken to reach the band’s upper (lower) edge. Thus, the maximum growth during the amplification is approximately given by
$N_{k}\simeq\exp{\left(\int_{-t_{\rm amp}/2}^{t_{\rm amp}/2}\frac{3\xi m_{P}^{2}H^{2}}{2f^{2}\dot{\theta}(t_{k})}dt\right)}$ (63)
$\simeq\exp{\left(\frac{(3\xi m_{P}^{2}H^{2})^{2}}{f^{4}\dot{\theta}(t_{k})^{2}|H(t_{k})\dot{\theta}(t_{k})+\ddot{\theta}(t_{k})|}\right)}.$ (64)
Hence, for the fluctuations to grow significantly, we need to have the term inside the exponent to be greater than unity. Using $\dot{\theta}(t)=\dot{\theta}_{m}\frac{H(t)}{H_{m}}$, $\ddot{\theta}(t)=-3H(t)\dot{\theta}(t)$ and our relation for $\dot{\theta}_{m}$ (see Eq. (21)), we arrive at the following bound on $\xi$
$\xi\lesssim\frac{4}{9}(12)^{3}\left(\frac{f}{m_{P}}\right)^{2}\simeq 768\left(\frac{f}{m_{P}}\right)^{2}\,.$ (65)
Comparing with Fig. 6, we can see that it agrees quite well with our numerical results [^13] This bound is even more relaxed because the growth of the fluctuations does not necessarily imply fragmentation of the axion, where the latter takes place only when the energy density of the fluctuations become comparable to that of the background homogeneous mode.. Therefore, the lower part of the range in Eq. (9) is quite safe against fragmentation of the rotating axion condensate, for which we can have successfull baryogenesis as well, as discussed in Section III. In fact, in the next section we argue that it may be preferable for $\xi$ to be not much larger than $(f/m_{P})^{2}$.
We leave a detailed study of axion fragmentation including improved analytical results and numerical solutions with backreaction effects in such a scenario for future studies.
@@ VIII The Kibble issue
Therefore, we expect the axion to roll only in one direction (e.g. clockwise) in the observable Universe, which evades the concerns of the Kibble problem. It also avoids another possible criticism of our setup. This has to do with the fact that the second-order phase transition which follows the change of tilt of the axion vacuum manifold, would be a waterfall transition, as in Ricci reheating, and is expected to proceed in a stochastic manner Bettoni et al. (2022); Figueroa and Byrnes (2017); Nakama and Yokoyama (2019), that might not end up with the dynamics we discussed in Sec, II. Having our axion displaced from the maximum of the effective potential, as suggested by Eq. (66), implies that the transition is not waterfall, and our treatment is better justified. The price to pay is that we have to make sure that excessive isocurvature perturbations are avoided. [^15] This is quite feasible for a rotating axion, e.g. see Ref. Chung and Tadepalli (2024).
Another important point is that $\xi$ has to gradually increase during kination if the axion is to roll at all, because the latter requires $V_{\rm eff}^{\prime\prime}||>H^{2}$. For this to happen $\xi$ needs to grow by a couple of orders of magnitude, if we are close to the upper bound in Eq. (67). One possibility is to consider a running $\xi(\sigma)$ as Bezrukov and Shaposhnikov (2014); Hamada et al. (2015); Ezquiaga et al. (2018); Drees and Xu (2021); Cheong et al. (2021); Ghoshal et al. (2024)
@@
i.e. near the edge of both ranges in Eqs. (9) and (67). It is such a value that was assumed in the example at the end of Sec. IV, while, as shown in Sec. VII, it is also safe from fragmentation of the axion condensate.
Recall that in this low range of $\xi$, our estimate of $\dot{\theta}_{m}$ given by Eq. (21) is unreliable. However, the relevant physical quantity to track is the average value of $\dot{\theta}$, which remains approximately equal to the potential barrier height (cf. Fig. 2). This is what we used in Eq. (21). Hence, our estimate of the baryon asymmetry is expected to remain valid as an order of magnitude, even with $\xi$ around the value given by Eq. (69). We leave a detailed numerical simulation for future studies.
Now, using the above value of $\xi$, we can find a lower bound on the RHN mass scale as follows. Combining Eqs. (26) and (69), it is straightforward to find
@@
where we used $Y_{B}\sim 10^{-10}$. Thus, if for example $T_{\rm reh}\sim 10^{10}\,$GeV we find $T_{\rm B-L}\sim 10^{9}\,$GeV. Considering the BBN bound $T_{\rm reh}\gtrsim 10^{7}$ GeV, we get $T_{\rm B-L}\gtrsim 3\times 10^{7}$ GeV. In the case of our concrete example, discussed in Sec. VI, for $z_{\rm fo}$ around $\mathcal{O}(10)$, this gives a lower bound on the RHN mass as $M_{N}\gtrsim 3\times 10^{8}\,$GeV.
@@ IX Conclusion
In this work, we investigated an attractive way to generate rotation for an axion-like particle (which is not the QCD axion), if it is non-minimally coupled to gravity, respecting the shift-symmetry. Such a rotation is generated when the effective potential of the axion flips at the end of inflation, providing the axion enough kinetic energy to surpass the potential barrier. This becomes possible if inflation is followed by a kination-dominated epoch ($w=1$), when the axion potential barrier redshifts in the same way as its kinetic energy, i.e. as $a^{-6}$. We find that for the axion to roll and remain sub-dominant, the value of $\xi$ is constrained to be in the range given by Eq. (9). We utilize such an axion rotation to generate the baryon asymmetry through spontaneous baryogenesis. The final baryon asymmetry, given by Eq. (26), is found to depend on the initial kinetic energy of the axion determined crucially by the non-minimal coupling $\xi$ and the decay constant $f$, and its subsequent redshift which depends on the reheating temperature $T_{\rm reh}$. Additionally, the asymmetry is determined by the decoupling temperature of the baryon or lepton number violating interaction $T_{\rm B-L}$, which depends on the details of the particle physics setup. The axion continues to rotate even after reheating until the kinetic energy becomes comparable to the height of the bare mass potential (given in Eq. (3)), determined by $M$, after which the coherent oscillation of the axion can behave as dark matter, relating the axion decay constant to its mass through Eq. (35).
The presence of the kination-dominated epoch leads to a substantial enhancement of the stochastic gravitational wave background from inflation. The blue-tilt of such a GW spectrum is determined by the reheating temperature. In order not to overproduce GWs during BBN, the reheating temperature is constrained to be greater than $\mathcal{O}(10^{7})\,$GeV (Eq. (49)), which substantially constrains the viable parameter space.
@@
We applied our mechanism to a concrete example of Majoron in the well-motivated Type I seesaw model, where our framework can be utilized for the cogenesis of baryon asymmetry and dark matter, with the viable parameter space shown in Fig. 5. It is important to note that in our mechanism of flipped axion rotation, although the same axion field is responsible for cogenesis, the potential (shown in Eq. (7)) playing role in baryogenesis during rotation differs from that (shown in Eq. (3)) generating the dark matter abundance during oscillation. Thus, the axion mass $m_{\phi}$ is unconstrained from the observed baryon asymmetry (cf. Eq. (26)), which opens up a larger parameter space unlike in other axion rotation scenarios. In our example, the Majoron mass is generally predicted to take any value in the range below sub-eVs. On the other hand, the decoupling temperature $T_{\rm B-L}$ is tightly constrained by bounds from GW overproduction and issues related to fragmentation and the Kibble problem discussed earlier. As a result, in our example, the RHN mass which determines $T_{\rm B-L}$ is constrained to be larger than around $10^{8}$ GeV (see the discussion after Eq. (70)). While considering $m_{\nu}\sim 0.05$ eV and the Yukawa couplings to obey the perturbativity bound, $M_{N}$ and $T_{\rm max}$ can be upto the GUT-scale.
Our scenario is very predictive and favors specific values of the non-minimal coupling $\xi$, which for instance, considering GUT-scale decay constant $f\sim 10^{-2}\leavevmode\nobreak\ m_{P}$ is around $10^{-4}$. These predictions can be indirectly tested by GW experiments like BBO, UDECIGO and resonant cavities, which can be sensitive to the blue tilt (see Fig. 4). Morever, the predicted values of $\Delta N_{\rm eff}$ coming from the GW background can be within the sensitivities of several experiments such as CMB-S4, CMB-HD, EUCLID (cf. Table 1, Figs. 4 and 5). In the example we considered, one can also have complementary tests in neutrino experiments (cf. Eq.  (54)). Our idea can also have interesting phenomenology for other particle physics models including the QCD axion, which we leave for future exploration.
@@ Acknowledgements
CC is supported by National Natural Science Foundation of China (No. 12433002) and Start-up Funds for Doctoral Talents of Jiangsu University of Science and Technology. SJD is supported by IBS under the project code IBS-R018-D1. KD is supported (in part) by the STFC consolidated Grant: ST/X000621/1. Authors thank Pawel Kozow for discussion. CC thanks the supports from The Asia Pacific Center for Theoretical Physics, The Center for Theoretical Physics of the Universe at Institute for Basic Science during his visits. The authors also thank Maximilian Berbig for some useful comments.